Shift report 20030202 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 02 2003 - 23:59:50 EST

  • Next message: Apache: "Shift report 20030203 00:00-08:00"
    ------------------------------------------------------------
    Date: Sun Feb  2 23:59:49 2003, Shift: 16:00-24:00
    Supervisor: JH Lee, with: 
    Log book#13 pages 142 - 142
    
    Shift Summary:
    ------------------------------------------------------------
      Beam:
      -40 min. beam ended with "Controls Failure"
      during steering at ~19:00.
      -No useful beam since then.
    
      Data:
      -Runs 7746-7751
      -MRS at 90 A350: 80K -> Done MRS 90 A350 300K
      -FS at 8 1/4A: 20K -> 120K so far. 180K more to go
      -MRS moved to 40deg. A500 -> Goal: 400K
    
      Problems:
      -T1 Very unstable: anode tripping
      Now at park mode.
    
      Plan for the next shift:
      -MRS at 40 A500: 400K
      -FS at 8 1/4A: 180K
      -Take data without T1 (keep T1 off)
      -Watch TOFW/ZDC at 40 degree
       (good beam: TOFW/ZDC ~20? at 40deg.: see the elog)
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 03 2003 - 00:00:43 EST